OpenStudy (e.mccormick):

Just checking. \[\vec{u}=(-2,-4,-3), \vec{v}=(-1,-3,-4); Proj_v\vec{u}\]

OpenStudy (e.mccormick):

\[\frac{\langle \vec{u},\vec{v} \rangle}{\langle \vec{v},\vec{v} \rangle}\vec{v}=\frac{\langle (-2,-4,-3),(-1,-3,-4) \rangle}{\langle (-1,-3,-4),(-1,-3,-4) \rangle}\vec{v}\\ \implies \frac{2+12+12}{1+9+16}\vec{v}\implies \frac{26}{26}\vec{v}\implies(-1,-3,-4)\]Which just seems too good to be true.

OpenStudy (dan815):

isnt that just the dot product xD

OpenStudy (e.mccormick):

Project \(\vec{u}\) onto \(\vec{v}\).

OpenStudy (e.mccormick):

The wonders of notation... they seem to nver agree on a single notation in Linear. No, the dot would be an orthogonal check.

OpenStudy (dan815):

or instead of that just do U.(V/|V|) where you are taking the unit direction vector of V that way U.Unit vector = U cos@ or component of U along V

OpenStudy (e.mccormick):

Projection is an orthornormal basis for a subspace...

OpenStudy (dan815):

what does that mean? so projection is the component of U on V

OpenStudy (dan815):